OpenStudy (anonymous):

The domain of the function 1/√16x+15 ?

OpenStudy (shubhamsrg):

is it (16x)^1/2 ?

OpenStudy (anonymous):

no its 1/ √(16x+15)

sam (.sam.):

\[x \in \mathbb{R} : x >- \frac{15}{16}\]

sam (.sam.):

Question is\[\frac{1}{\sqrt{16 x+15}}\]

OpenStudy (shubhamsrg):

gotcha!

OpenStudy (anonymous):

can u put it in interval?